OpenStudy (anonymous):

complex numbers: find a, b such that (-bi)^2 = a^2 ?

OpenStudy (anonymous):

a^2+b^2=0; a=m+ni b=p+qi (m+ni)^2+(p+qi)^2=0 m^2+p^2=n^2+q^2....(1) mn=pq;....(2) now choose m,n,p,q such that the conditions 1,2 satisfied simultaneously....

OpenStudy (anonymous):

are a and b complex numbers ?

OpenStudy (anonymous):

yes... its given..

OpenStudy (anonymous):

how did you get mn = pq

OpenStudy (anonymous):

sorry that will be mn=-pq.

OpenStudy (anonymous):

also you foiled wrong

OpenStudy (anonymous):

(m+ni)^2+(p+qi)^2 = m^2 + 2mni -n^2 + p^2 + 2pqi - q^2 = 0

OpenStudy (anonymous):

so???

OpenStudy (anonymous):

where is 2mni and 2pqi

OpenStudy (anonymous):

2mn=2pq => mn=pq

OpenStudy (anonymous):

in your algebra , i dont see it

OpenStudy (anonymous):

i have done it in single step.

OpenStudy (anonymous):

too fast for me

OpenStudy (anonymous):

sorry im doing another problem, helping student, one sec

OpenStudy (anonymous):

i dont follow your logic

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

2mn i = -2pqi

OpenStudy (anonymous):

ok then mn = -pq, agreed

OpenStudy (anonymous):

something tells me that a,b are integers though originally

OpenStudy (anonymous):

btw: a and b are real

OpenStudy (anonymous):

the correct answer is: a=b?

OpenStudy (anonymous):

our guy here thinks that a and b are complex numbers , i thought this was simpler, lol

OpenStudy (anonymous):

the question is: find a and b (both real) such that (-bi)^2 = a^2

OpenStudy (anonymous):

if a, b are not complex then the relation (a)^2=(-bi)^2 does not hold.

OpenStudy (anonymous):

so a=b=0

OpenStudy (anonymous):

right

OpenStudy (anonymous):

see your question starts with "complex numbers". so i thought that

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

this equation only have a trivial solution..

OpenStudy (anonymous):

yeah, no fun

OpenStudy (anonymous):

why a = b = 0? is it because a^2 + b^2 = 0 + 0i?

OpenStudy (anonymous):

so the question is more interesting if you allow a and b to be complex , right?

OpenStudy (anonymous):

for real no. only a=b=0 holds

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a^2 = (-b*i)^2 = b^2*i^2 = b^2 *-1 ,

OpenStudy (anonymous):

can you find any such a,b other than 0 that holds your condition????????????????

OpenStudy (anonymous):

so a^2 = -b^2 , so a^2 + b^2 = 0,

OpenStudy (anonymous):

no because a^2>=0 , and b^2 >= 0

OpenStudy (anonymous):

see i think a,b must be considered as complex otherwise its very easy....

OpenStudy (anonymous):

just solve the damn equation, lol

OpenStudy (anonymous):

if you take them complex then also you will find a solution 0+0i along with some new solution...

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

so this equation is equilvalent, find all a,b, such that a^2 + b^2 = 0, for a,b element of C

OpenStudy (anonymous):

damn...i have another typo: -(bi)^2 = a^2...therefore b^2 = a^2 - so the solution is a = b?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

noooooooooooooooooooooooooooooooooo

OpenStudy (anonymous):

(-bi)^2 does not mean -(bi)^2

OpenStudy (anonymous):

yes i know

OpenStudy (anonymous):

find a and b (both real) such that -(bi)^2 = a^2.......this is the correct problem statement.....sorry....